Hi. I am really struggling with this one.

A rectangular piece of cardboard whose length is twice its width is used to construct an open box. Cutting a 1 foot by 1 foot square off of each corner and folding up the edges will yield an open box. If the desired volume is 12 cubic feet, what are the dimensions of the original rectangular piece of cardboard??
Thank you!

Answer by scott8148(5880)

let 2x=length, so x=width __ each dimension is reduced by 2 when the corners are cut

(2x-2)(x-2)(1)=12 __ 2x^2-6x+4=12

subtracting 12 and dividing by 2 __ x^2-3x-4=0

factoring __ (x-4)(x+1)=0